• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:02:08 on February20,2009

    Entry number 262335

    keyword=Shift summary (day)

    Shift started with continuation of runs 20285 and 1623 (HRS mom 1.42)
    8:48. Run 1623 is over. Changing momentum to 1.270 GeV. Stopped beam to put the N2 ref cell in.

    8:55 Left HRS momentum ready, 5uA on Reference Cell N2 with 102psi on display,
    LHRS run 20806 8:55-9:34 Q=0.0115
    BB run 1625 (bigbite_buf_fadc mode) 9:04-9:34 Q=0.0091

    (BB run 1624 had too high deadtime)

    9:35 Move to polarized 3He cell, change BB DAQ configuration to bigbite_buf;

    9:42 - taking data on LHRS. There is problem with BB buffer mode (cannot write to data file) and Brad is fixing it;

    10:00 Target NMR 63.43%

    10:15 BB DAQ problem fixed, start taking data on BB.

    13:30 Hall in controlled access to fix network speed.

    14:10 Target NMR 63.43%

    14:30-15:00 follow beam checkout procedure with 2uA.
    Run 20290 and 1634 were spot check with optics target, raster 5.5x6

    15:00-16:00 Production with 15 uA.
    BB run 1635-1642 were deadtime check.


    Left HRS pol3He runs at Kinematics H:

    20287 09:42-10:00, 10uA, 67K events, Q=0.007695
    20288 10:06-12:07, 10uA, 500K events, Q=0.05981
    20289 12:08-13:25, 10uA, 326K events, Q=0.03983
    20291 14:57-swing shift, 15uA

    Bigbite pol3He runs at Kinematics H:

    1628 10:14-12:02, 7M events, Q=0.05389
    1629 12:04-13:25, 5.6M events, Q=0.04179
    1643 15:39-swing shift, 15uA